Trying to figure out the answer to: (x-2)^1/3=2
Answer by lwsshak3(11628)

(x-2)^1/3=2
take log of both sides
(1/3)log(x-2)=log2
log(x-2)=3log2=log(2^3)
x-2=2^3=8
x=10
check:
(x-2)^1/3
=(10-2)^(1/3)
=8^(1/3)
=2